OpenStudy (anonymous):

An investor bought 3 shares of stock a and 2 shares of stock b for a total of $41. Stock a costs $2.00 more per share than stock b. What is the cost of a share of stock a in dollars?

OpenStudy (campbell_st):

well you know how much you spent 3a + 2b = 41 and you also know that when comparing prices a = b + 2 so you can substitute the 2nd equation into the 1st equation and eliminate a 3(b + 2) + 2b = 41 or 3b + 6 + 2b = 41 which simplifies to 5b + 6 = 41 all you need to do is solve for b. when you have b substitute it into a = b + 2 to find a. hope this helps.